# The work was supported by the grant for research centers in the field of AI provided by the Ministry of Economic Development of the Russian Federation in accordance with the agreement  [^00] C313925P4F0002 and the agreement with Skoltech №139-10-2025-033

### v4

@@ II Related Work
For personalized models, AriGraph [anokhin2024arigraphlearningknowledgegraph] introduces a framework that integrates episodic memory and long-term planning using knowledge graphs. Similarly, HippoRAG [gutiérrez2024hipporagneurobiologicallyinspiredlongterm] employs personalized algorithms to improve question-answering (QA) performance by constructing semantic graphs, demonstrating notable advancements over conventional extraction methods. MemWalker [chen2023walkingmemorymazecontext] and PEARL [sarthi2024raptorrecursiveabstractiveprocessing] address challenges associated with context length, proposing architectures capable to efficiently traversing and consolidating information from large-scale documents.
@@ VI Experiments and Results
Retrieval algorithm 7B 8B 14B+ worse configs (w restr) best configs (w restr) other configs (w/o restr) worse configs (w restr) best configs (w restr) other configs (w/o restr) worse configs (w restr) best configs (w restr) other configs (w/o restr) WC 0.09 0.34 0.55 AS 0.1 0.175 0.14 0.29 0,36 0.41 0.23 0.36 0.33 BS 0.025 0.18 0.06 0.26 0.5 0.36 0.48 0.6 0.65 WC+BS 0.033 0.095 0.02 0.30 0.39 0.32 0.62 0.7 0.68 BS+AS 0.02 0.175 0.01 0.25 0.48 0.36 0.48 0.64 0.66 AS+WC 0.055 0.115 0.07 0.33 0.37 0.42 0.57 0.6 0.6
TABLE IV: Stability of proposed retrieval algorithms when various restrictions are imposed: ”worse–configs (w restr)” – low quality configurations and imposed restrictions; ”best–configs (w restr)” – high quality configurations and imposed restrictions; ”other–configs (w/o restr)” – configurations without restrictions on graph traversal
@@ VIII Future work
In future work, we propose to enhance the temporal dynamics of our memory graph by introducing a ”memory time” parameter, which will enable fine-grained filtering of triples based on temporal proximity and edge types. This modification will allow the system to selectively prioritize temporally proximate data or emphasize specific relationship categories, thereby improving the precision of personalized responses. Also, one possible future direction is to explore erasure-coded and locally recoverable layouts for sharding graph and vector indices across nodes, inspired by information-theoretic distributed storage, enabling fast repair and continued operation under partial server unavailability [10.1134/S1064226920120116]. In addition, we will investigate private and verifiable retrieval protocols (PIR with result verification) so an agent can query remote memory without revealing the user’s intent and can detect incorrect or malicious responses [10.48550/arXiv.2301.11730]. These mechanisms aim to make long-term personalized memory robust, secure, and auditable at scale.

### v5

# arXiv:2506.17001v5[cs.CL] 27 Feb 2026

@@ II Related Work
For personalized models, AriGraph [anokhin2024arigraphlearningknowledgegraph] introduces a framework that integrates episodic memory and long-term planning using knowledge graphs. Similarly, HippoRAG [gutiérrez2024hipporagneurobiologicallyinspiredlongterm] employs personalized algorithms to improve question-answering (QA) performance by constructing semantic graphs, demonstrating notable advancements over conventional extraction methods. MemWalker [chen2023walkingmemorymazecontext] and RAPTOR [sarthi2024raptorrecursiveabstractiveprocessing] address challenges associated with context length, proposing architectures capable to efficiently traversing and consolidating information from large-scale documents.
@@ VI Experiments and Results
Retrieval algorithm 7B 8B 14B+ worse configs (w restr) best configs (w restr) other configs (w/o restr) worse configs (w restr) best configs (w restr) other configs (w/o restr) worse configs (w restr) best configs (w restr) other configs (w/o restr) WC 0.09 0.34 0.55 AS 0.1 0.175 0.14 0.29 0.36 0.41 0.23 0.36 0.33 BS 0.025 0.18 0.06 0.26 0.5 0.36 0.48 0.6 0.65 WC+BS 0.033 0.095 0.02 0.30 0.39 0.32 0.62 0.7 0.68 BS+AS 0.02 0.175 0.01 0.25 0.48 0.36 0.48 0.64 0.66 AS+WC 0.055 0.115 0.07 0.33 0.37 0.42 0.57 0.6 0.6 Mean: 0.046 0.148 0.065 0.286 0.42 0.36 0.47 0.58 0.57
TABLE IV: Stability of proposed retrieval algorithms when various restrictions are imposed: ”worse–configs (w restr)” – low LLM-as-a-Judge score configurations and imposed restrictions; ”best–configs (w restr)” – high LLM-as-a-Judge score configurations and imposed restrictions; ”other–configs (w/o restr)” – configurations without restrictions on graph traversal
@@ VIII Future work
In future work, we, firstly, propose to enhance the temporal dynamics of our memory graph by introducing a ”memory time” parameter, which will enable fine-grained filtering of triples based on temporal proximity and edge types. This modification will allow the system to selectively prioritize temporally proximate data or emphasize specific relationship categories, thereby improving the precision of personalized responses. Secondly, recognizing potential bottlenecks associated with current implementation of graph traversal algoritms, our efforts also will be focused on fine-tuning underlying vector storage schemes employed alongside advanced approximate nearest neighbor search techniques. These enhancements promise substantial reductions in overall query latencies while maintaining comparable precision rates. Thirdly, to reduce vector search space and speed up vector retrieval operations we plan to add more characteristics by which triples from the knowledge graph can be aggregate and stored in a separate smaller, but more concentrated and specific vector stores.
Also, one possible future direction is to explore erasure-coded and locally recoverable layouts for sharding graph and vector indices across nodes, inspired by information-theoretic distributed storage, enabling fast repair and continued operation under partial server unavailability [10.1134/S1064226920120116]. In addition, we will investigate private and verifiable retrieval protocols (PIR with result verification) so an agent can query remote memory without revealing the user’s intent and can detect incorrect or malicious responses [10.48550/arXiv.2301.11730]. These mechanisms aim to make long-term personalized memory robust, secure, and auditable at scale.